OpenStudy (anonymous):

find the derivative function of x^2 e^tan(x) find the derivative function of x^2 e^tan(x) @Mathematics

OpenStudy (anonymous):

you have to use the product rule

OpenStudy (anonymous):

is the answer x^2 (sex^2(x) e^tan(x) ) + e^tan(x) (2x) ?

OpenStudy (anonymous):

x^2 (sec^2(x) e^tan(x) ) + e^tan(x) (2x) ?

OpenStudy (anonymous):

use \[(fg)'=f'g+g'f\] with \[f(x)=x^2,f'(x)=2x, g(x)=e^{\tan(x)}, g'(x)=\sec^2(x)e^{\tan(x)}\]